OpenStudy (this):

10x^2-x-2=0 a, -240;two real solution b, 21;two imaginary solution c, 81;two real solution d, -3;two imaginary solution

OpenStudy (this):

some one help me plzz

OpenStudy (tkhunny):

We'll need to know the question. What is asked? Solve for 'x'? Produce of the zeros? Characterize the zeros?

OpenStudy (tkhunny):

Discriminant: (-1)^2 -4(10)(-2) = 1 + 80 = 81 - Two Real

OpenStudy (this):

helpppp